please help me solve this equation: if 3x-1=11, what is the value of 2x+x?
Thank you all in advance!!!!!!

Answer by funmath(2933)

if 3x-1=11, what is the value of 2x+x?
First, solve the equation:
3x-1=11
3x-1+1=11+1
3x=12
3x/3=12/3
x=4
Now substitute x=4 into 2x+x
2(4)+(4)
8+4
12
Happy Calculating!!!